OpenStudy (anonymous):

1. f(x) = x/x^2 + 1 2. e^2x + e^-x a) Find the intervals on which f is increasing or decreasing b) find local max and min c) find intervals of concavity and inflection points Thank You!

OpenStudy (anonymous):

\[f(x)=\frac{x}{x^2+1}\] \[f'(x)=\frac{1-x^2}{(x^2+1)^2}\] denominator is always positive, so this is positive if the numerator is

OpenStudy (anonymous):

solve \(1-x^2>0\) since this is a parabola that faces down, it is positive between the zeros, namely on the interval \((-1,1)\) and negative otherwise

OpenStudy (anonymous):

therefore function is decreasing on \((-\infty,-1)\cup (1,\infty)\) and increasing on \((-1,1)\) as you can clearly see from the picture here http://www.wolframalpha.com/input/?i=x%2F%28x^2%2B1%29